OpenStudy (anonymous):

Line AB contains points A (1, 2) and B (−2, 6). The slope of line AB is zero undefined positive negative

OpenStudy (anonymous):

slope=y2-y1/x2-x1 plug in your points 6-2/-2-1 =4/-3 =-4/3 so your answer is "negative"

OpenStudy (anonymous):

Thank you! That is what I got the first time but I wasn't sure.